OpenStudy (anonymous):

What are the x- and y-intercepts of the line? 2x – 3y = 12 A. x-intercept –6, y-intercept 4 B. x-intercept –4, y-intercept 6 C. x-intercept 4, y-intercept –6 D. x-intercept 6, y-intercept –4

OpenStudy (anonymous):

I tried every single one and I cant find it.

OpenStudy (***[isuru]***):

hi, divide the whole equation from 3 2x – 3y = 12 u will get... (2/3)x - y = 4 y = (2/3)x -4 now this equation is in the form of y = mx + b where b is the y intercept... so.. in our equation y intercept is " -4" to find x intercept... u have to find the value of "x" when "y" is 0... y = (2/3)x -4 0 = (2/3)x - 4 (2/3)x = 4 2x = 12 x = 6 so.. the x intercept is "6" hope this will help ya!!!

OpenStudy (anonymous):

ohh ok thnx

OpenStudy (***[isuru]***):

u r welcome!!

OpenStudy (anonymous):

y= 2x-12 ----- 3 plug in 0 for x and y = -4 (0,-4) x= 12+3y ------ 2 plug in y=0 to get x=6 (6,0)

OpenStudy (anonymous):

How is it 6,0 I thought @***[ISURU]*** said D (6,-4)